OpenStudy (anonymous):

If f(x) = 2x2 - x, find f(-3). 15 21 39

OpenStudy (anonymous):

f(-3) is just saying: let x = -3

OpenStudy (anonymous):

So, f(-3) = 2(-3)^2 + 3

OpenStudy (anonymous):

f(-3) = 2(9) + 3 f(-3) = 18 + 3 f(-3) = 21

OpenStudy (anonymous):

It's like this, when they say 'Find f(-3)', they mean "let x = -3".

OpenStudy (anonymous):

So whenever you see 'x' in your equation, you replace it with -3.

OpenStudy (anonymous):

So if I had: f(x) = 2x + 4 And it said: "Find f(1)" I would have: f(1) = 2(1) + 4